Write the following numbers in scientific notation:

1. 125.365
2. 0.45
3. 5000
1
Expert's answer
2020-09-09T17:20:33-0400

5000 = 5*103

0.45 = 4.5 * 10-1

125.365 = 1.25365 * 105